For the matrix A-1-4-3 2 3) consider the set 2 S-b E Rb-Ax for some xE R The vector -4 E S since -4Ax where X- The vector -3 E Ssince -3 Ay where The vector 7 E S since -7 Az where
Answer #1

It may be obseved that (3,-4,2)T = 1(3,-4,2)T+0(3,-3,3)T. Hence x = (1,0)T.

Also, (3,-3,3)T = 0(3,-4,2)T+0(3,-3,3)T . Hence y = (0,1)T

and (6,-7,5)T =1(3,-4,2)T1(3,-3,3)T. Hence z = (1,1)T